OpenStudy (anonymous):

What are the solutions of x^2 + 8x − 20 = 0? Show all work for full credit.

OpenStudy (phi):

To solve x^2 + 8x − 20 = 0 I would factor it to find the factors: first, notice the -20, that means the two factors will be opposite signs second, the +8 (on the x term) means the larger factor is plus third, list the factors of 20 1,20 2,10 4,5 does any pair when subtracted give us +8? if so, that is the pair to use : (x-2)(x+10)= 0 now solve for x: either x-2=0 or x+10=0

OpenStudy (anonymous):

x^2+10x-2x-20=0 x(x+10)-2(x+10)=0 (x+10)(x-2)=0 either or x+10=0 x-2=0 x=-10 x=2